OpenStudy (anonymous):

write the standard form equation of each hyperbola: vertices: (1,9), (1, -15) endpoints of conjugate axis: (14, -3), (-12,-3)

OpenStudy (dumbcow):

vertices along line x=1 implying vertical hyperbola 2a = |-15-9| = 24 a = 12 a^2 = 144 conjugate axis along y=-3 2b = |-12-14| = 26 b=13 b^2 = 169 center (1,-3) \[\frac{(y+3)^{2}}{144} - \frac{(x-1)^{2}}{169} = 1\]